# Enhanced Curvature Perturbation and Primordial Black Hole Formation in Two-stage Inflation with a break

### v2

@@ 1 Introduction
There are a number of proposed mechanisms to enhance the primordial curvature perturbations during inflation, for instance, single-field inflation with a non-slow-roll stage  [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80], multi-field inflation [81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102], modify gravity [103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132], curvaton scenario [133, 134, 135, 136, 137, 138], parametric resonance [139, 140, 141, 142], etc. The property that any successful model must have is its consistency with the CMB and LSS constraints on large scales. In other words, it is essential to decouple the physics on small scales from that on large scales. This can be naturally realized in the scenario where inflation consists of two or more distinct periods intervened by a transition stage that may or may not break inflation. We call it the two-stage inflation. It was proposed by various people for various purposes [143, 144, 145, 146, 147, 148].
@@ 2 The model
$S_{J}=\int d^{4}x\sqrt{-g}\ \bigg{[}\frac{M_{\mathrm{pl}}^{2}}{2}f(R)-\frac{1}{2}g^{\mu\nu}\partial_{\mu}\chi\partial_{\nu}\chi-V(\chi)\bigg{]},$ (2.1)
where $M_{\mathrm{pl}}\equiv(8\pi G)^{-1/2}$, $f(R)$ is given by
$f(R)=R+\frac{R^{2}}{6M^{2}}-\frac{\xi R}{M_{\mathrm{pl}}^{2}}(\chi-\chi_{0})^{2},$ (2.2)
@@
to guarantee vanishing of the cosmological constant at the end of inflation.
@@
the action (2.1) can be expressed in the Einstein frame as [150, 151, 152, 153, 154, 155]
@@
    After the scalaron $\phi$ rolls down from the potential plateau, it undergoes damped oscillations around its local minimum at $\phi_{c}$, which is a function of $\chi$, and the first stage of inflation ends. In our scenario, as we consider the energy scale of the second inflationary stage to be much smaller than that of the first stage, it follows that $\phi_{c}\ll M_{\rm pl}$. Then due to the rapid decay of $\phi$ to $\phi_{c}$, $F(\phi)$ rapidly approaches unity, and the Hubble parameter $H$ drops down by several orders of magnitude. As the rapid change of $H$ means the slow-roll parameter $\epsilon\equiv-\dot{H}/H^{2}$ of order unity, inflation is temporarily halted until the potential of $\chi$ starts to dominate. During this stage, the effective mass square of $\chi$ changes from positive to negative, which makes $\chi$ start rolling and drive the second stage of inflation.
@@
Let us now comment on the role of $\chi_{0}$ in (2.2). As explained above, or as shown in Fig. 1, $\chi$ is trapped at the local minimum $\chi\approx\chi_{0}$ during St-1. If $\chi_{0}=0$, as soon as St-2 starts at which $\phi$ settles down to the local minimum $\phi=\phi_{c}$, the point $\chi=0$ becomes a local maximum, implying that classically $\chi$ would be stuck $\chi=0$, and it would move away from $\chi=0$ only by the quantum fluctuations. This leads to unacceptably large curvature perturbations. A non-vanishing $\chi_{0}$ avoids this catastrophe. The same strategy is also taken in [156] to avoid the overproduction of topological defects which often happens in hybrid inflation scenarios such as [82, 86]. Before closing this section, let us briefly comment on the stability of the trajectory during St-1. During most of this stage, we have $F\gg 1$ so that the effective mass of $\chi$ is positive and $\chi$ acts as a damped harmonic oscillator (See Fig. 1). In our model, we find that the coupling parameter $\xi$ must satisfy $\xi>3/16$ to ensure that $\chi$ undergoes an under-damped oscillators during St-1 so that it settles down to the stable local minimum $\chi_{0}$ in a few e-folds. This secures that St-1 can be effectively treated as single-field slow-roll inflation.
@@ 3 The Background Solution
Refer to captionRefer to caption
Figure 3: The background solution. The left panel shows the numerical solution for the background equations. The right panel shows the evolution of the slow-roll parameter $\epsilon_{H}$. It is obvious that during stage 2, the slow-roll condition is violated, hence resulting in a break between two stages of inflation.
@@ 3.2 St-2
where we have introduce the parameter $A$ defined by
@@
Figure 4: The numerical result of the effective mass of the two scalar fields. The blue, red solid line and the black dotted line shows the time evolution of effective mass of $\phi$, effective mass of $\chi$, and $H^{2}$. The parameters used for this plot can be found in case 1 of Table. 1. The two vertical solid lines indicates from the left to right the horizon exiting time for modes $k_{2}$ and $k_{1}$.
@@
we have used (3.5) to express $V_{0}$ and $M^{2}$ in terms of $\mu^{2}$. Thus total e-folding number of St-2 is estimated as
@@ 3.3 St-3
$+\mu^{-2}\left[1-\xi\left(\frac{\chi-\chi_{0}}{M_{\mathrm{pl}}}\right)^{2}\right]^{-1}\left[1+\frac{1}{2}\left(\frac{\chi}{\chi_{g}}\right)^{2}+\frac{1}{4}\left(\frac{\chi}{\chi_{g}}\right)^{4}\right],$ (3.34)
where $\chi_{g}=m/\sqrt{\lambda}$ is the value of $\chi$ at the global minimum, firstly mentioned below (2.3). It is evident that $\phi_{c}$ evolves along with the evolution of $\chi$, as illustrated in the lower right panel of Fig. 1. Similar to the way we derived (3.26), we set $\phi(n)=\phi_{c}(n)+\Delta\phi(n)$ with $\Delta\phi(n)\ll\phi_{c}(n)$ during St-3. Inserting this ansatz into (2.12), neglecting the contribution of $\xi(\chi-\chi_{0})^{2}/M_{\mathrm{pl}}^{2}$ and $\chi/\chi_{g}$ terms in the regime where $\chi\ll\chi_{g}\ll M_{\rm pl}$, the equation of motion for $\Delta\phi(n)$ can be approximated as
@@ 4.1 Preliminaries
where $\delta\phi$ and $\delta\chi$ are those defined on flat slicing.
@@ 4.3 Superhorizon evolution and $\delta N$
Actually this result can be immediately derived from the expression for $N_{1}$ given by (3.12). The number of e-folds is independent of $\chi$ due to the single-field slow-roll nature of St-1.
@@ 5.1 Primordial black hole formation
$\beta(M)=\frac{\gamma}{\sqrt{2\pi\sigma_{\delta}^{2}(M_{H})}}\int_{\delta_{\mathrm{th}}}^{\infty}\exp\left(-\frac{\delta^{2}}{2\sigma_{\delta}^{2}(M_{H})}\right)\mathrm{d}\delta=\gamma\leavevmode\nobreak\ \mathrm{erfc}\left(\frac{\delta_{\mathrm{th}}}{\sqrt{2}\sigma_{\delta}(M_{H})}\right),$ (5.2)
where $\sigma_{\delta}^{2}(M_{H})$ is the variance of the density contrast on scale $M_{H}$, smoothed by a window function $W(k;R)$ to exclude the effect of small scale perturbations,
@@
$\mathcal{P}_{\delta}(k)=\left(\frac{4}{9}\right)^{2}(kR)^{4}P_{\mathcal{R}}(k)\,.$ (5.4)
@@ 7 Acknowledgement
We thank Shi Pi, Jinsu Kim, and Minxi He for useful discussions. This work is supported in part by JSPS KAKENHI Grant Nos. JP20H05853 and JP24K00624. XW is supported by China Scholarship Council No. 202306260140. YZ is supported by the Fundamental Research Funds for the Central Universities, and by Project 12047503 supported by NSFC. XW and YZ gratefully acknowledge the hospitality and support of the Kavli Institute of Physics and Mathematics of the Universe (Kavli IPMU), the University of Tokyo.
@@ Appendix A A general description of the field equations
where $\dot{\psi}^{a}\equiv d\psi^{a}/dt$, while ${D\dot{\psi}^{a}}/{dt}\equiv\dot{\psi}^{a}+\Gamma^{a}_{bc}\dot{\psi}^{b}\dot{\psi}^{c}$ is the covariant derivative in the curved field space.
@@
Appendix B Parameters
@@
We summarize the constraints that the parameters should satisfy in order for our model to be viable in Table. 2.

### v3

@@ 1 Introduction
There are a number of proposed mechanisms to enhance the primordial curvature perturbations during inflation, for instance, single-field inflation with a non-slow-roll stage  [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83], multi-field inflation [84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105], modify gravity [106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135], curvaton scenario [136, 137, 138, 139, 140, 141], parametric resonance [142, 143, 144, 145], etc. The property that any successful model must have is its consistency with the CMB and LSS constraints on large scales. In other words, it is essential to decouple the physics on small scales from that on large scales. This can be naturally realized in the scenario where inflation consists of two or more distinct periods intervened by a transition stage that may or may not break inflation. We call it the two-stage inflation. It was proposed by various people for various purposes [146, 147, 148, 149, 150, 151].
@@ 2 The model
$S_{J}=\int d^{4}x\sqrt{-g}\ \bigg{[}\frac{M_{\mathrm{pl}}^{2}}{2}f(R,\chi)-\frac{1}{2}g^{\mu\nu}\partial_{\mu}\chi\partial_{\nu}\chi-V(\chi)\bigg{]},$ (2.1)
where $M_{\mathrm{pl}}\equiv(8\pi G)^{-1/2}$, $f(R,\chi)$ is given by
$f(R,\chi)=R+\frac{R^{2}}{6M^{2}}-\frac{\xi R}{M_{\mathrm{pl}}^{2}}(\chi-\chi_{0})^{2},$ (2.2)
@@
to guarantee the vanishing of the cosmological constant at the end of inflation.
@@
the action (2.1) can be expressed in the Einstein frame as [153, 154, 155, 156, 157, 158, 159]
@@
    After the scalaron $\phi$ rolls down from the potential plateau, it undergoes damped oscillations around its local minimum at $\phi_{c}$, which is a function of $\chi$, and the first stage of inflation ends. In our scenario, as we consider the energy scale of the second inflationary stage to be much smaller than that of the first stage, it follows that $\phi_{c}\ll M_{\rm pl}$. Then due to the rapid decay of $\phi$ to $\phi_{c}$, $F(\phi)$ rapidly approaches unity, and the Hubble parameter $H$ drops down by several orders of magnitude. As the rapid change of $H$ means the slow-roll parameter $\epsilon\equiv-\dot{H}/H^{2}$ of order unity, inflation is temporarily halted until the potential of $\chi$ starts to dominate. During this stage, the effective mass square of $\chi$ changes from positive to negative, which makes $\chi$ start rolling and drives the second stage of inflation.
@@
Let us now comment on the role of $\chi_{0}$ in (2.2). As explained above, or as shown in Fig. 1, $\chi$ is trapped at the local minimum $\chi\approx\chi_{0}$ during St-1. If $\chi_{0}=0$, as soon as St-2 starts at which $\phi$ settles down to the local minimum $\phi=\phi_{c}$, the point $\chi=0$ becomes a local maximum, implying that classically $\chi$ would be stuck $\chi=0$, and it would move away from $\chi=0$ only by the quantum fluctuations. This leads to unacceptably large curvature perturbations. A non-vanishing $\chi_{0}$ avoids this catastrophe. The same strategy is also taken in [160] to avoid the overproduction of topological defects which often happens in hybrid inflation scenarios such as [85, 89]. Before closing this section, let us briefly comment on the stability of the trajectory during St-1. During most of this stage, we have $F\gg 1$ so that the effective mass of $\chi$ is positive and $\chi$ acts as a damped harmonic oscillator (See Fig. 1). In our model, we find that the coupling parameter $\xi$ must satisfy $\xi>3/16$ to ensure that $\chi$ undergoes an under-damped oscillator during St-1 so that it settles down to the stable local minimum $\chi_{0}$ in a few e-folds. This secures that St-1 can be effectively treated as single-field slow-roll inflation.
@@ 3 The Background Solution
Figure 3: The background solution. The left panel shows the numerical solution for the background equations. The right panel shows the evolution of the slow-roll parameter $\epsilon_{H}$. It is obvious that during stage 2, the slow-roll condition is violated, hence resulting in a break between the two stages of inflation.
@@ 3.2 St-2
where we have introduced, the parameter $A$ defined by
@@
Figure 4: The numerical result of the effective mass of the two scalar fields. The blue, red solid line and the black dotted line show the time evolution of the effective mass of $\phi$, the effective mass of $\chi$, and $H^{2}$. The parameters used for this plot can be found in case 1 of Table. 1. The two vertical solid lines indicate from the left to right the horizon exiting time for modes $k_{2}$ and $k_{1}$.
@@
we have used (3.5) to express $V_{0}$ and $M^{2}$ in terms of $\mu^{2}$. Thus the total e-folding number of St-2 is estimated as
@@ 3.3 St-3
$+\mu^{-2}\left[1-\xi\left(\frac{\chi-\chi_{0}}{M_{\mathrm{pl}}}\right)^{2}\right]^{-1}\left[1-\left(\frac{\chi}{\chi_{g}}\right)^{2}\right]^{2},$ (3.34)
where $\chi_{g}\equiv m/\sqrt{\lambda}$ is the value of $\chi$ at the global minimum, firstly mentioned below (2.3). It is evident that $\phi_{c}$ evolves along with the evolution of $\chi$, as illustrated in the lower right panel of Fig. 1. Similar to the way we derived (3.26), we set $\phi(n)=\phi_{c}(n)+\Delta\phi(n)$ with $\Delta\phi(n)\ll\phi_{c}(n)$ during St-3. Inserting this ansatz into (2.12), neglecting the contribution of $\xi(\chi-\chi_{0})^{2}/M_{\mathrm{pl}}^{2}$ and $\chi/\chi_{g}$ terms in the regime where $\chi\ll\chi_{g}\ll M_{\rm pl}$, the equation of motion for $\Delta\phi(n)$ can be approximated as
@@ 4.1 Preliminaries
where $\delta\phi$ and $\delta\chi$ are those defined on the flat slicing. The detailed derivation of these expressions is given in Appendix B.
@@ 4.3 Superhorizon evolution and $\delta N$
This result can be immediately derived from the expression for $N_{1}$ given by (3.12). The number of e-folds is independent of $\chi$ due to the single-field slow-roll nature of St-1.
@@ 5.1 Primordial black hole formation
$\beta(M)=\frac{\gamma}{\sqrt{2\pi\sigma_{\delta}^{2}(H)}}\int_{\delta_{\mathrm{th}}}^{\infty}\exp\left(-\frac{\delta^{2}}{2\sigma_{\delta}^{2}(H)}\right)\mathrm{d}\delta=\gamma\leavevmode\nobreak\ \mathrm{erfc}\left(\frac{\delta_{\mathrm{th}}}{\sqrt{2}\sigma_{\delta}(H)}\right),$ (5.2)
where $\sigma_{\delta}^{2}(H)$ is the variance of the density contrast on scale $H$, smoothed by a window function $W(k;R)$ to exclude the effect of small scale perturbations,
@@
$\mathcal{P}_{\delta}(k)=\left(\frac{4}{9}\right)^{2}(kR)^{4}\mathcal{P}_{\mathcal{R}}(k)\,.$ (5.4)
@@ 7 Acknowledgement
We thank Shi Pi, Jinsu Kim, and Minxi He for useful discussions. This work is supported in part by JSPS KAKENHI Grant Nos. JP20H05853 and JP24K00624. XW is supported by China Scholarship Council No. 202306260140, YZ is supported by the Fundamental Research Funds for the Central Universities, and by Project 12047503 supported by NSFC, and by China Scholarship Council. XW and YZ gratefully acknowledge the hospitality and support of the Kavli Institute of Physics and Mathematics of the Universe (Kavli IPMU), the University of Tokyo.
@@ Appendix A A general description of the field equations
where $\dot{\psi}^{a}\equiv d\psi^{a}/dt$, while ${D\dot{\psi}^{a}}/{dt}\equiv\frac{d\dot{\psi}^{a}}{dt}+\Gamma^{a}_{bc}\dot{\psi}^{b}\dot{\psi}^{c}$ is the covariant derivative in the curved field space.
@@
Appendix B The comoving curvature and isocurvature perturbations
In this appendix, following [91, 121], we show the explicit expressions for the comoving curvature perturbation and the isocurvature perturbation in our model.
The comoving slicing can be considered for each field, on which the field value is homogeneous. The curvature perturbation on the comoving slices for the field $\psi^{a}$ is defined as
$\mathcal{R}_{c}^{a}\equiv\mathcal{R}+\frac{H}{\dot{\psi}^{a}}\delta\psi^{a}=\frac{H}{\dot{\psi}^{a}}Q^{a},$ (B.1)
where $\mathcal{R}$ and $\delta\psi^{a}$ are the curvature perturbation and the field fluctuations evaluated on an arbitrary slicing, and $Q^{a}$ are the field fluctuations on the flat slicing $\mathcal{R}=0$. Then, the total comoving curvature perturbation is defined as
$\mathcal{R}_{c}\equiv\mathcal{R}+H\frac{h_{ab}\dot{\psi}^{a}\delta\psi^{b}}{\dot{\psi}_{0}^{2}}\,,$ (B.2)
where
$\dot{\psi}_{0}^{2}\equiv h_{ab}\dot{\psi}^{a}\dot{\psi}^{b}=\dot{\phi}^{2}+F^{-1}\dot{\chi}^{2}.$ (B.3)
To evaluate the comoving curvature perturbation and the isocurvature perturbation, we work in the flat slicing $\mathcal{R}=0$. The field perturbation can be decomposed into the components parallel and orthogonal to the background trajectory,
$Q_{\parallel}$ $\equiv h_{ab}T^{a}Q^{b}=\frac{\dot{\phi}\delta\phi+F^{-1}\dot{\chi}\delta\chi}{\sqrt{\dot{\phi}^{2}+F^{-1}\dot{\chi}^{2}}},$ (B.4)
$Q_{\perp}$ $\equiv h_{ab}N^{a}Q^{b}\,,$ (B.5)
where $T^{a}$ and $N^{a}$ are the unit vectors parallel and normal to the trajectory, respectively, given by
$T^{a}$ $\equiv\frac{\dot{\psi}^{a}}{|\dot{\psi}_{0}|}=\frac{1}{\sqrt{\dot{\phi}^{2}+F^{-1}\dot{\chi}^{2}}}\left(\dot{\phi},\dot{\chi}\right),$ (B.6)
$N^{a}$ $\equiv\frac{D{T^{a}}/dt}{\left|D{T^{a}}/dt\right|}\,;\quad\frac{DT^{a}}{dt}=\frac{dT^{a}}{dt}+\Gamma^{a}_{bc}\dot{\psi}^{b}T^{c}\,.$ (B.7)
In our model, we have
$\frac{DT^{1}}{dt}$ $=\frac{1}{\dot{\psi}_{0}^{3}}F^{-1}\dot{\chi}\left({\frac{D\dot{\phi}}{dt}\dot{\chi}-\frac{D\dot{\chi}}{dt}\dot{\phi}}\right)\,,$ (B.8)
and similarly
$\frac{DT^{2}}{dt}=\frac{1}{\dot{\psi}_{0}^{3}}F^{-1/2}\dot{\phi}\left(\frac{D\dot{\chi}}{dt}\dot{\phi}-\frac{D\dot{\phi}}{dt}\dot{\chi}\right).$ (B.9)
Inserting (B.8) and (B.9) into (B.7), we obtain
$N^{a}=\frac{F^{1/2}}{\dot{\psi}_{0}}(F^{-1}\dot{\chi},-\dot{\phi})\,.$ (B.10)
The curvature perturbation on the comoving slice $\mathcal{R}_{c}$ and the isocurvature perturbation $\mathcal{S}$ are given by
$\mathcal{R}_{c}$ $=\frac{H}{\dot{\psi}_{0}}Q_{\parallel}\,,$ (B.11)
$\mathcal{S}$ $=\frac{H}{\dot{\psi}_{0}}Q_{\perp}\,.$ (B.12)
Note that (B.11) is the same as (B.2). Inserting (B.10) into (B.5), together with (B.4), we obtain the explicit expressions for $\mathcal{R}_{c}$ and $\mathcal{S}$ as
$\mathcal{R}_{c}$ $=H\frac{\dot{\phi}\delta\phi+F^{-1}\dot{\chi}\delta\chi}{\dot{\phi}^{2}+F^{-1}\dot{\chi}^{2}}\,,$ (B.13)
$\mathcal{S}$ $=HF^{-1/2}\frac{\dot{\chi}\delta\phi-\dot{\phi}\delta\chi}{\dot{\phi}^{2}+F^{-1}\dot{\chi}^{2}}\,.$ (B.14)
Appendix C Parameters
@@
We summarize the constraints that the parameters should satisfy for our model to be viable in Table. 2.